must show all work and calculations.
Answer #1


apply snells law na sinA = nb sin B

where A is angle of incidence, B is angle of Refraction

na and nb are refractive index

a. nb = 1* sin 41.5/sin (25.4)

nb = 1.544

b. for criticla, angle B = 90 deg

so sin A = nb/na = 1/1.544

sinA = 0.647

A = 40.36 deg
Answer #2

n1*sini1 = n2*sinr1

1*sin41.5 = n2*sin25.4

n2= 1.54

a) glass

b) sinC = 1/n2

C = 40.49